Of 5000 vehicles sold by an auto dealership last month, 200 have a defect that
will require a factory recall for dealer correction. For a simple random sample
of n = 60 of last month’s customers, what is the probability that no more than
20 will be affected by the recall?
Answer #1

Ans.

P(defect) =200/5000 = 0.04

P(X<=20) = P(X=0)+.....P(X=20) = 1.000